40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 746     Entry time: Mon Jul 28 11:20:13 2008
Author: Jenne 
Type: Update 
Category: PSL 
Subject: Work on the FSS and Reference Cavity 
[Yoichi, Jenne, Koji]

The Reference Cavity's saga continues....

Thursday, Yoichi and I worked to change the beam that we chose from the 2nd pass through the AOM, to the first order beam rather than the 2nd order beam (see elog #726). After choosing the correct beam, we get 29mW incident on the reference cavity (compared with 4mW before any work began). We adjusted the angle of the AOM in the plane of the table, and got up to 30.6mW. We adjusted the tip/tilt of the AOM and got to 30.7mW (the tip/tilt adjustment made a more significant difference in the work described in elog #726, but after that work, it was probably already pretty close to optimized). We noticed that for the above measurements, we had 2 beams through the Polarizing Beam Splitter and Waveplate (one very dim), so after excluding that beam, the power meter read 30.4mW. We adjusted the curved mirror a little, and got 30.8mW incident on the reference cavity.

We then put a triangle wave into the offset of the MC Servo Board using the "trianglewave <channel> <center> <amplitude> <period> <runtime>" command in a terminal screen. This changes the voltage to the VCO, and thus the frequency response of the AOM. We watch the diffracted spots from the second pass through the AOM, and confirm that the beam we have chosen is not moving, and all the others are. By symmetry, if we chose the first order beam after the first pass through the AOM, and then again chose the first order beam after the second pass, the resulting beam will not move with the frequency change of the AOM.

We saw 1.50V (Refl. PD, unlocked) on the 'scope after aligning the optics to make the newly chosen beam hit the input mirror of the reference cavity. Order of operations for this alignment:
  • Recenter the beam on the 2 lenses that are just after the PBS and the waveplate
  • Adjust pitch and yaw of the two steering mirrors until the beam reflected off the input mirror of the reference cavity is parallel to the incident beam
    • Use a sensor card to check the alignment of the incident and reflected beams, and adjust the steering mirrors to get the alignment close
      • Note the amplitude of the DC output of the Refl. PD with the iris completely open. Close the iris until the signal decreases by ~50%, then adjust the steering mirrors until the original amplitude is regained. Repeat until the iris can be almost completely closed but the Refl. PD signal doesn't change
    • Watch the DC output of the Refl. PD, and maximize the signal on a 'scope
    • Sweep the PZT of the laser using a function generator into the RAMP input on the FSS board (~10Vpp at ~1Hz), OR sweep the temperature of the laser using the trianglewave function on the SLOW FSS channel (amplitude~0.5, period~50)
    • Watch the modes that resonate in the cavity, and adjust pitch and yaw of the steering mirrors to get closer to the TEM00 mode
    • When the TEM00 mode appears in the sweep, stop the sweep, and lock the cavity
    • Watch the DC output of the Transmitted PD, and maximize the signal on a 'scope
  • Celebrate!

After all of this adjusting,
Refl. PD (unlocked) = 1.48V
Refl. PD (locked) = 680mV
Trans. PD (locked) = 6.28V
Power reflected (unlocked) = 26.28mW
Power transmitted (locked) = 13.89mW
Thus, 53% transmission

Next: check the amount of power transmitted by reducing the amplitude of the RF modulator. This reduces the amount of power used by the sidebands, and so should increase the transmission.
Power incident = 27mW
Power transmitted = 17.2mW
Thus, 64% transmission
We then put the RF modulator back where it was originally.

We then replaced the lens mounts for the f=802 and f=687 lenses between the AOM and the reference cavity, to the new mounts that Yoichi bought. Koji helped me realign into the reference cavity, and we got:
Refl. PD (unlocked) = 1.48V
Refl. PD (locked) = 880mV
Trans PD (locked) = 4.64V
Power incident = 26.97mW
Power transmitted = 10.39mW
39% transmission
Since more mode matching etc. is in the works, we left this for the night.

On Friday, we changed the setup of the cameras and PDs for both reflection and transmission, to avoid saturating the PDs and cameras.

On the Refl. side of the reference cavity, we put a W2-PW-1025-UV-1064-45P pickoff between the last mirror and lens before the camera and PD. We moved the camera to the pickoff side of the new optic. We then replaced teh 45UNP beam splitter that split the beam between the PD and the camera with a Y1-1037-45P highly reflective mirror, and put the PD in the old camera location.

On the Trans. side of the ref. cavity, we replaced the BSI-1064-50-1025-45S with a W2 pickoff, and replaced the Y1-1037-45-P highly reflective mirror with the 50/50 beam splitter that was replaced by the W2.

Now we have:
Refl. PD (unlocked) = 1.68V
Refl. PD (locked) = 640mV
Trans PD (locked) = 4.24V
Power incident = 25mW
Power transmitted = 14.48mW
58% transmission

Koji pointed out that when remounting, I had put the f=802 lens ~2cm away from its original position (along the z-axis), so I moved the lens back to where it should be, and realigned into the reference cavity. Since Rana was working on the PMC at the same time, the laser was turned down by about a factor of 100, so my starting measurements were:
Refl. PD (unlocked) = 23.6mV
Refl. PD (locked) = 10.2mV
Trans PD (locked) = 56mV
Power incident = 0.35mW
Power transmitted = 0.16mW
46% transmission

Since it was late on Friday by the time everything was realigned into the ref. cavity (I'm still working on my optics aligning skills), I forgot to measure the transmission after all of my work. I'll do that today (Monday) as soon as Sharon/Koji are done working with the IFO this morning. Also, I'll put up before/after pictures as soon as I find the camera...it seems to have walked off.

UPDATE:
Ref. Cav. measurements after Friday's alignment (and after turning the laser power back up to normal):
Refl. PD (unlocked) = 1.58V
Refl. PD (locked) = 304mV
Trans PD (locked) = 3.68V
Power incident = 24.96mW
Power transmitted = 16.45mW
66% transmission


To do: Start the actual mode-matching into the reference cavity.
ELOG V3.1.3-